Given the two reactions PbCl2⇌Pb2++2Cl−, K3 = 1.72×10−10, and AgCl⇌Ag++Cl−, K4 = 1.29×10−4, what is the equilibrium constant Kfinal for the following reaction? PbCl2+2Ag+⇌2AgCl+Pb2+
Answer #1

PbCl2⇌Pb2++2Cl−, K3 = 1.72×10−10

K3    = [Pb^2+][Cl^-]^2/[PbCl2]

AgCl⇌Ag++Cl−, K4 = 1.29×10−4

K4   = [Ag^+][Cl^-]/[AgCl]

K4^2    = [Ag^+]^2[Cl^-]^2/[AgCl]^2

K3*K4^2   = [Pb^2+][Cl^-]^2/[PbCl2] * [Ag^+]^2[Cl^-]^2/[AgCl]^2

                 = [AgCl]^2[Pb^2+]/[PbCl2][Ag^+]^2

PbCl2+2Ag+⇌2AgCl+Pb2+

Kc   = [AgCl]^2[Pb^2+]/[PbCl2][Ag^+]^2

kc    = K3*K4^2

        = 1.72*10^-10 * (1.29*10^-4)^2

        = 2.86*10^-18

the equilibrium constant K final =2.86*10^-18 >>>>>answer